# An Explicit M1 Radiation-hydrodynamics Scheme for Three-dimensional Protostellar Evolution

### v1

@@ Abstract
We present a radiation-hydrodynamics (RHD) scheme that enables three-dimensional simulations resolving both protostellar interiors and their surrounding accretion flows within a single framework, to clarify how a protostar evolves while interacting with the accretion flow. The method builds on an explicit M1 closure scheme with a reduced speed of light approximation (RSLA) for massively parallel computation. Our scheme introduces a complementary non-RSLA radiation component that dominates in optically thick regions. This hybrid treatment restores physical energy conservation inside protostars, which would otherwise be violated under the RSLA, while retaining the advantage of large time steps. To overcome the limitation of the conventional M1 closure in solving radiative transfer in extremely optically thick regions inside protostars and across steep optical–depth gradients near their surfaces, we incorporate the optical–depth information of neighboring cells into the radiative transfer calculation. We further evolve photon number densities in addition to radiation energy densities to reconstruct an effective local spectrum on the fly without resorting to costly multi-frequency transport. We implement this scheme in the adaptive mesh refinement code SFUMATO and verify its validity through a series of test calculations. As an application, we follow the early evolution of a massive protostar formed at high redshift, within a full cosmological context. The results reveal a continuous structure connecting the swollen protostar and its surrounding disk, which cannot be captured in conventional one-dimensional models. This explicit RHD scheme opens a path to studies of protostellar evolution and its interaction with the accretion flow in realistic three-dimensional environments.
@@ 1 INTRODUCTION
The protostellar evolution with mass accretion represents the fundamental stage during which a star gains most of its mass. The nature of this process determines the final stellar masses and, in turn, the shape of the stellar mass spectrum across cosmic time, both in the early universe and in present-day star-forming environments (see recent reviews by R. S. Klessen & S. C. O. Glover, 2023; P. Hennebelle & M. Y. Grudić, 2024; H. Beuther et al., 2025).
@@
Protostars evolve through complex interactions with the accreting gas surrounding them. The accretion flow supplies entropy and angular momentum to the protostars, thereby influencing their internal structure and evolutionary path. Conversely, massive protostars exert strong radiative and mechanical feedback on the surrounding gas, regulating the subsequent accretion process. In recent studies, large-scale accretion dynamics under the protostellar feedback have primarily been investigated with three-dimensional (3D) radiation hydrodynamic (RHD) simulations (e.g. M. R. Krumholz et al., 2007a; R. Kuiper et al., 2011; H. Susa, 2013; M. Klassen et al., 2016; A. L. Rosen et al., 2016; T. Hosokawa et al., 2016; A. Stacy et al., 2016; S. Chon et al., 2018; K. Sugimura et al., 2020; O. Jaura et al., 2022; K. Sugimura et al., 2023; S. Chon et al., 2024; P. Sharda & S. H. Menon, 2025). In such simulations, however, the protostar itself is typically treated with sub-grid models, such as sink particles, and neither its internal structure nor the boundary layer between the protostar and the accretion disk is resolved explicitly. Meanwhile, the structure and evolution of protostars on sub-au scales have been studied with one-dimensional (1D) stellar evolution calculations (e.g. S. W. Stahler et al., 1980, 1986; F. Palla & S. W. Stahler, 1991; R. Behrend & A. Maeder, 2001; K. Omukai & F. Palla, 2001, 2003; H. W. Yorke & P. Bodenheimer, 2008; T. Hosokawa & K. Omukai, 2009; T. Hosokawa et al., 2010, 2013; Y. Sakurai et al., 2015; L. Haemmerlé et al., 2018a, b; N. P. Herrington et al., 2023; D. Nandal et al., 2023, 2024). These 1D models typically assume spherically symmetric accretion onto the protostar, a hydrostatic structure within the stellar interior, and a simple rotational motion of the protostar itself, thereby neglecting both the inherently anisotropic nature of gas inflow and the resulting complex dynamics within the protostar. This separation of scales has limited our ability to fully capture the mutual influence between the protostar and its accreting environment.
@@
To overcome this limitation, it is essential to resolve both the stellar interior and the large-scale accretion flow in a unified 3D RHD simulation. Recent developments in computational resources and numerical schemes have made such simulations increasingly feasible (Y. Luo et al., 2018; A. Bhandare et al., 2020, 2024, 2025; K. Kimura et al., 2023; A. Ahmad et al., 2023, 2024; A. C. Mayer et al., 2025). A widely adopted radiative transfer (RT) method in these simulations is the flux-limited diffusion (FLD) approximation with implicit time integration. The FLD method reduces the computational cost by replacing the full RT problem, which involves three spatial, two angular, and one frequency dimensions, with a single zeroth moment diffusion equation (C. D. Levermore & G. C. Pomraning, 1981). Moreover, the implicit time integration enables us to take time steps longer than the timescale required for stable explicit integration. This is particularly advantageous in star formation simulations, where the dynamical timescale of the gas is much longer than the RT timescale.
@@
In addition to the FLD method with implicit time integration, the M1 closure method combined with the reduced speed of light approximation (RSLA) has also been widely employed in modern star formation RHD simulations (C. D. Levermore, 1984; M. A. Skinner & E. C. Ostriker, 2013; J. Rosdahl & R. Teyssier, 2015; R. Kannan et al., 2019; R. Mignon-Risse et al., 2020; H. Fukushima & H. Yajima, 2021; A. Yaghoobi et al., 2025). The M1 closure scheme solves not only the zeroth but also the first moment of the RT equation, which allows for a more accurate treatment of anisotropic radiation fields and prevents unphysical diffusion compared to the FLD. In addition, the RSLA alleviates the severe time step constraints imposed by the Courant–Friedrichs–Lewy (CFL) condition in M1 closure, enabling stable explicit integration with large time steps (N. Y. Gnedin & T. Abel, 2001). More sophisticated approaches, such as the Variable Eddington Tensor method, have been used to overcome the limitations inherent to the M1 closure scheme, for instance its inability to accurately treat multi-beam radiation fields (e.g., J. M. Stone et al., 1992; P. Sharda & S. H. Menon, 2025). However, owing to the balance between accuracy and computational cost, the M1 closure scheme still remains widely adopted in current large-scale star formation simulations.
@@
Such an explicit formulation is highly suitable for massively parallel computing and GPU-accelerated architectures, and can outperform FLD with implicit time integration depending on the simulation setup. Therefore, to follow realistic protostellar evolution over longer timescales in massively parallel simulations, it is necessary to apply the M1 closure scheme with the RSLA to simulations that resolve both the protostellar interior and the large-scale accretion flow simultaneously. However, direct application of the M1 closure scheme with the RSLA gives rise to the following challenges.
@@
The first challenge is ensuring total energy conservation in protostars. When the RSLA is employed, the rate of energy exchange between gas and radiation is artificially reduced to reproduce the steady-state of the system with the full speed of light. However, this modification causes the loss of physical energy during the energy exchange, thereby leading to violation of total energy conservation. This issue is not problematic in optically thin regions outside protostars, where energy naturally escapes from the system as radiation. In optically thick environments, such as the interiors of protostars, however, this artificial energy loss effectively removes entropy from the system, preventing accurate modeling of protostellar evolution.
@@
The second challenge concerns the treatment of RT within the protostellar interior and across its surface. In star formation RHD simulations, the computational grid is typically refined according to the Jeans criterion, i.e., resolving the Jeans length by a prescribed number of cells. However, even when the Jeans length is resolved with several to a few tens of cells, the optical depth per grid cell in the protostellar interior still far exceeds unity. In such regions, the M1 closure method introduces excessive numerical diffusion to ensure integration stability. To mitigate this issue, J. Rosdahl & R. Teyssier (2015) proposed a scheme that splits the radiation energy into two components, which effectively suppresses artificial diffusion and reproduces the physical behavior of radiative diffusion in optically thick interiors (see also C.-C. He et al., 2024). At the protostellar surface, however, steep gradients of optical depth naturally arise, where radiation transitions from being trapped to freely escaping. In this regime, their method of splitting the radiation energy into two components prevents radiation from readily leaking across the sharp gradient, causing the scheme to underestimate the outgoing radiation flux despite its accuracy in the interior.
@@
The third challenge concerns the treatment of the radiation energy spectrum, which is a common limitation of previous RHD simulations including those using the FLD method. Spectral information is essential for accurately evaluating the protostellar radiative feedback on the surrounding gas, including heating and ionization. However, many RHD simulations assume a fixed spectral shape rather than following its evolution explicitly. For example, it is common to adopt a spectrum based on simplified 1D protostellar models from previous studies. However, in simulations resolving the protostar itself that we focus on here, the spectral evolution may deviate significantly from these simplified models. Thus, we need to estimate it on-the-fly in a manner consistent with the actual protostellar evolution in simulations. The most straightforward approach to determining the radiation spectrum is to divide the energy into multiple frequency bins and solve the emission, absorption, and transport processes for each bin (H. W. Yorke & C. Sonnhalter, 2002; R. Kuiper et al., 2010; M. Klassen et al., 2014; A. L. Rosen et al., 2017; R. Kuiper & T. Hosokawa, 2018). While this method provides high accuracy, it substantially increases computational cost. Therefore, an alternative method with lower computational cost is required.
@@
In this study, we propose a new RHD scheme that addresses the challenges discussed above. Our RT approach is based on the explicit M1 closure method with the RSLA, which follows the anisotropy of the radiation field and ensures high parallel efficiency for large-scale computations. We have implemented the scheme in the self-gravitating magnetohydrodynamics code SFUMATO (T. Matsumoto, 2007; T. Matsumoto et al., 2015), which has been widely applied to various problems of star formation (e.g. T. Inoue et al., 2018; T. Matsumoto et al., 2019; K. Sugimura et al., 2020; H. Fukushima et al., 2020; H. Fukushima & H. Yajima, 2021; D. Abe et al., 2021; K. E. Sadanari et al., 2021; K. Kimura et al., 2021, 2023; K. Sugimura et al., 2023; K. E. Sadanari et al., 2024). Through a series of tests, we demonstrate the accuracy and robustness of the method. Furthermore, we apply this scheme to primordial star formation in the early universe and show how protostars evolve in conjunction with the surrounding large-scale gas dynamics.
@@
The remainder of this paper is organized as follows. In Section 2, we review the basic RHD formulation of the M1 closure scheme, combined with the multicomponent method proposed by J. Rosdahl & R. Teyssier (2015) and the RSLA. In Section 3, we provide an overview of the three challenges discussed above. Section 4 describes the formulation of our improved RHD scheme to solve the challenges. In Section 5, we present its numerical implementation in the self-gravitating magnetohydrodynamics code, SFUMATO. In Section 6, we validate our method using a series of test problems. Section 7 presents simulation results for the evolution of a protostar formed in the early universe as an application of our new scheme. Finally, Section 8 provides a summary of this paper.
2 Review of RHD Formulation
@@
This section presents the RHD formulation of the M1 closure scheme with the multicomponent method proposed by J. Rosdahl & R. Teyssier (2015) and the RSLA. In Section 2.1, we introduce the basic RHD equations for star formation simulations, where the gas velocity is much smaller than the speed of light. In Section 2.2, we describe the method by J. Rosdahl & R. Teyssier (2015), which modifies the M1 closure scheme to follow RT accurately in highly optically thick regions. In Section 2.3, we review the RSLA formulation, which relaxes the CFL constraint imposed by the speed of light and allows for larger time steps.
2.1 Basic RHD equations with M1 closure
@@
Here, we introduce the RHD equations. Throughout this paper, we restrict our discussion to the regime where $\beta=|\bm{v}|/c\ll 1$, with $\bm{v}$ denoting the gas velocity and $c$ the speed of light. We take into account both absorption and scattering processes, including their frequency dependence. The formulation of the RHD equations depends on the choice of reference frame, specifically, whether physical quantities are expressed in the lab frame or in the comoving frame. In this study, we adopt the so-called mixed frame formulation, in which opacities and emissivities are evaluated in the comoving frame, while all other gas and radiation quantities are defined in the lab frame (D. Mihalas & R. I. Klein, 1982; M. R. Krumholz et al., 2007b). This formulation enables the numerical algorithm to conserve total energy to machine precision during gas–radiation interaction. Moreover, because radiation energy in the lab frame is a conserved quantity, it is obvious how to refine or coarsen the grid in a conservative manner in AMR codes. In this paper, physical quantities measured in the comoving frame are denoted with the subscript 0. We show the Lorentz transformations between the lab frame and comoving frame quantities in Appendix A.
@@
$\kappa_{\mathrm{0E}}=\frac{\int_{0}^{\infty}d\nu\hskip 2.84526pt\kappa^{\nu}_{0}E^{\nu}_{\mathrm{rad},0}}{\int_{0}^{\infty}d\nu\hskip 2.84526ptE^{\nu}_{\mathrm{rad},0}},$ (9)
$\chi_{\mathrm{0F}}=\frac{\int_{0}^{\infty}d\nu\hskip 2.84526pt(\kappa^{\nu}_{0}+\sigma^{\nu}_{0})F^{\nu}_{0}}{\int_{0}^{\infty}d\nu\hskip 2.84526ptF^{\nu}_{0}},$ (10)
where $\kappa^{\nu}_{0}$ and $\sigma^{\nu}_{0}$ are the frequency-dependent absorption and scattering opacities evaluated in the comoving frame, $B^{\nu}(T)$ is the Planck distribution at temperature $T$, and $E^{\nu}_{\mathrm{rad},0}$ and $F^{\nu}_{0}$ are the comoving frame radiation energy density and flux per unit frequency. For calculating $\kappa_{0E}$ and $\chi_{0F}$ with Equations (9) and (10), we need the spectrum of radiation energy $E^{\nu}_{\mathrm{rad},0}$ and flux $F^{\nu}_{0}$. We describe how to calculate this spectrum with finite frequency bins and low computational cost in Section 4.3.
In the M1 closure method, we determine the radiation pressure tensor $\bm{\mathsf{P}}_{\mathrm{rad}}$ by imposing the M1 closure relation. Specifically, $\bm{\mathsf{P}}_{\mathrm{rad}}$ is expressed in terms of the Eddington tensor $\bm{\mathsf{D}}$ as follows (C. D. Levermore, 1984):
@@
With these definitions, the M1 closure reproduces the correct physical behavior of radiation in both the optically thin and optically thick limits. In the optically thin limit where $|F|/(cE_{\mathrm{rad}})\simeq 1$, $\xi\simeq 1$, the Eddington tensor $\bm{\mathsf{D}}$ approaches $\bm{n}\otimes\bm{n}$, corresponding to a beam-like radiation field. In the optically thick limit where $|F|/(cE_{\mathrm{rad}})\simeq 0$, $\xi\simeq 1/3$, it approaches $\bm{\mathsf{I}}/3$, consistent with an isotropic radiation field.
2.2 Streaming/trapped component for an optically thick medium

### v2

@@ Abstract
We present a radiation-hydrodynamics (RHD) scheme that enables three-dimensional simulations resolving both protostellar interiors and their surrounding accretion flows within a single framework, to clarify how a protostar evolves while interacting with the accretion flow. The method builds on an explicit two-moment M1 closure scheme with a reduced speed of light approximation (RSLA) for massively parallel computation. Our scheme introduces a complementary non-RSLA radiation component that dominates in optically thick regions. This hybrid treatment restores physical energy conservation inside protostars, which would otherwise be violated under the RSLA, while retaining the advantage of large time steps. To overcome the limitation of the conventional M1 closure in solving radiative transfer in extremely optically thick regions inside protostars and across steep optical-depth gradients near their surfaces, we incorporate the optical-depth information of neighboring cells into the radiative transfer calculation. We further evolve photon number densities in addition to radiation energy densities to reconstruct an effective local spectrum on the fly without resorting to costly multi-frequency transport. We implement this scheme in the adaptive mesh refinement code SFUMATO and verify its validity through a series of test calculations. As an application, we follow the early evolution of a massive protostar formed at high redshift, within a full cosmological context. The results reveal a continuous structure connecting the swollen protostar and its surrounding disk, which cannot be captured in conventional one-dimensional models. This RHD scheme opens a path to studies of protostellar evolution and its interaction with the accretion flow in realistic three-dimensional environments.
@@ I INTRODUCTION
The protostellar evolution with mass accretion represents the fundamental stage during which a star gains most of its mass. The nature of this process determines the final stellar masses and, in turn, the shape of the stellar mass spectrum across cosmic time, both in the early universe and in present-day star-forming environments (see recent reviews by Klessen and Glover, 2023; Hennebelle and Grudić, 2024; Beuther et al., 2025).
@@
Protostars evolve through complex interactions with the accreting gas surrounding them. The accretion flow supplies entropy and angular momentum to the protostars, thereby influencing their internal structure and evolutionary path. Conversely, massive protostars exert strong radiative and mechanical feedback on the surrounding gas, regulating the subsequent accretion process. In recent studies, large-scale accretion dynamics under the protostellar feedback have primarily been investigated with three-dimensional (3D) radiation hydrodynamic (RHD) simulations (e.g. Krumholz et al., 2007b; Kuiper et al., 2011; Susa, 2013; Klassen et al., 2016; Rosen et al., 2016; Hosokawa et al., 2016; Stacy et al., 2016; Chon et al., 2018; Sugimura et al., 2020; Jaura et al., 2022; Sugimura et al., 2023; Chon et al., 2024; Sharda and Menon, 2025). In such simulations, however, the protostar itself is typically treated with sub-grid models, such as sink particles, and neither its internal structure nor the boundary layer between the protostar and the accretion disk is resolved explicitly. Meanwhile, the structure and evolution of protostars on sub-au scales have been studied with one-dimensional (1D) stellar evolution calculations (e.g. Stahler et al., 1980, 1986; Palla and Stahler, 1991; Behrend and Maeder, 2001; Omukai and Palla, 2001, 2003; Yorke and Bodenheimer, 2008; Hosokawa and Omukai, 2009; Hosokawa et al., 2010, 2013; Sakurai et al., 2015; Haemmerlé et al., 2018a, b; Herrington et al., 2023; Nandal et al., 2023, 2024). These 1D models typically assume spherically symmetric accretion onto the protostar, a hydrostatic structure within the stellar interior, and a simple rotational motion of the protostar itself, thereby neglecting both the inherently anisotropic nature of gas inflow and the resulting complex dynamics within the protostar. This separation of scales has limited our ability to fully capture the mutual influence between the protostar and its accreting environment.
@@
To overcome this limitation, it is essential to resolve both the stellar interior and the large-scale accretion flow in a unified 3D RHD simulation. Recent developments in computational resources and numerical schemes have made such simulations increasingly feasible (Luo et al., 2018; Bhandare et al., 2020, 2024, 2025; Kimura et al., 2023; Ahmad et al., 2023, 2024; Mayer et al., 2025). A widely adopted radiative transfer (RT) method in these simulations is the flux-limited diffusion (FLD) approximation with implicit time integration. The FLD method reduces the computational cost by replacing the full RT problem, which involves three spatial, two angular, and one frequency dimensions, with a single zeroth moment diffusion equation (Levermore and Pomraning, 1981). Moreover, the implicit time integration enables us to take time steps longer than the timescale required for stable explicit integration. This is particularly advantageous in star formation simulations, where the dynamical timescale of the gas is much longer than the RT timescale.
@@
In addition to the FLD method with implicit time integration, the two-moment M1 closure method combined with the reduced speed of light approximation (RSLA) has also been widely employed in modern star formation RHD simulations (Levermore, 1984; Skinner and Ostriker, 2013; Rosdahl and Teyssier, 2015; Kannan et al., 2019; Mignon-Risse et al., 2020; Fukushima and Yajima, 2021; Yaghoobi et al., 2025). The M1 closure scheme solves not only the zeroth but also the first moment of the RT equation, which allows for a more accurate treatment of anisotropic radiation fields and prevents unphysical diffusion compared to the FLD. In addition, the RSLA alleviates the severe time step constraints imposed by the Courant–Friedrichs–Lewy (CFL) condition in M1 closure, enabling stable explicit integration with large time steps (Gnedin and Abel, 2001). More sophisticated approaches, such as the Variable Eddington Tensor method, have been used to overcome the limitations inherent to the M1 closure scheme, for instance its inability to accurately treat multi-beam radiation fields (e.g., Stone et al., 1992; Sharda and Menon, 2025). However, owing to the balance between accuracy and computational cost, the M1 closure scheme still remains widely adopted in current large-scale star formation simulations.
@@
Such an explicit formulation is highly suitable for massively parallel computing and GPU-accelerated architectures, and can outperform FLD with implicit time integration depending on the simulation setup (Wibking and Krumholz, 2022). Therefore, to follow realistic protostellar evolution over longer timescales in massively parallel simulations, it is necessary to apply the M1 closure scheme with the RSLA to simulations that resolve both the protostellar interior and the large-scale accretion flow simultaneously. However, direct application of the M1 closure scheme with the RSLA gives rise to the following challenges.
@@
The first challenge is ensuring total energy conservation in protostars. When the RSLA is employed, the rate of energy exchange between gas and radiation is artificially reduced to reproduce the steady-state of the system with the full speed of light. However, this modification causes the loss of physical energy during the energy exchange, thereby leading to violation of total energy conservation (Wibking and Krumholz, 2022). This issue is not problematic in optically thin regions outside protostars, where energy naturally escapes from the system as radiation. In optically thick environments, such as the interiors of protostars, however, this artificial energy loss effectively removes entropy from the system, preventing accurate modeling of protostellar evolution.
@@
The second challenge concerns the treatment of RT within the protostellar interior and across its surface. In star formation RHD simulations, the computational grid is typically refined according to the Jeans criterion, i.e., resolving the Jeans length by a prescribed number of cells. However, even when the Jeans length is resolved with several to a few tens of cells, the optical depth per grid cell in the protostellar interior still far exceeds unity. In such regions, the M1 closure method introduces excessive numerical diffusion to ensure integration stability. To mitigate this issue, Rosdahl and Teyssier (2015) proposed a scheme that splits the radiation energy into two components, which effectively suppresses artificial diffusion and reproduces the physical behavior of radiative diffusion in optically thick interiors (see also He et al., 2024). At the protostellar surface, however, steep gradients of optical depth naturally arise, where radiation transitions from being trapped to freely escaping. In this regime, their method of splitting the radiation energy into two components prevents radiation from readily leaking across the sharp gradient, causing the scheme to underestimate the outgoing radiation flux despite its accuracy in the interior.
@@
The third challenge concerns the treatment of the radiation energy spectrum, which is a common limitation of previous RHD simulations including those using the FLD method. Spectral information is essential for accurately evaluating the protostellar radiative feedback on the surrounding gas, including heating and ionization. However, many RHD simulations assume a fixed spectral shape rather than following its evolution explicitly. For example, it is common to adopt a spectrum based on simplified 1D protostellar models from previous studies. However, in simulations resolving the protostar itself that we focus on here, the spectral evolution may deviate significantly from these simplified models. Thus, we need to estimate it on-the-fly in a manner consistent with the actual protostellar evolution in simulations. The most straightforward approach to determining the radiation spectrum is to divide the energy into multiple frequency bins and solve the emission, absorption, and transport processes for each bin (Yorke and Sonnhalter, 2002; Kuiper et al., 2010; Klassen et al., 2014; Rosen et al., 2017; Kuiper and Hosokawa, 2018). While this method provides high accuracy, it substantially increases computational cost. Therefore, an alternative method with lower computational cost is required.
@@
In this study, we propose a new RHD scheme that addresses the challenges discussed above. Our RT approach is based on the explicit M1 closure method with the RSLA, which follows the anisotropy of the radiation field and ensures high parallel efficiency for large-scale computations. We have implemented the scheme in the self-gravitating magnetohydrodynamics code SFUMATO (Matsumoto, 2007; Matsumoto et al., 2015), which has been widely applied to various problems of star formation (e.g. Inoue et al., 2018; Matsumoto et al., 2019; Sugimura et al., 2020; Fukushima et al., 2020; Fukushima and Yajima, 2021; Abe et al., 2021; Sadanari et al., 2021; Kimura et al., 2021, 2023; Sugimura et al., 2023; Sadanari et al., 2024). Through a series of tests, we demonstrate the accuracy and robustness of the method. Furthermore, we apply this scheme to primordial star formation in the early universe and show how protostars evolve in conjunction with the surrounding large-scale gas dynamics.
@@
The remainder of this paper is organized as follows. In Section II, we review the basic RHD formulation of the M1 closure scheme, combined with the multicomponent method proposed by Rosdahl and Teyssier (2015) and the RSLA. In Section III, we provide an overview of the three challenges discussed above. Section IV describes the formulation of our improved RHD scheme to solve the challenges. In Section V, we present its numerical implementation in the self-gravitating magnetohydrodynamics code, SFUMATO. In Section VI, we validate our method using a series of test problems. Section VII presents simulation results for the evolution of a protostar formed in the early universe as an application of our new scheme. In Section VIII, we discuss the limitation of our new scheme. Finally, Section IX provides a summary of this paper.
II Review of RHD Formulation
@@
This section presents the RHD formulation of the M1 closure scheme with the multicomponent method proposed by Rosdahl and Teyssier (2015) and the RSLA. In Section II.1, we introduce the basic RHD equations for star formation simulations, where the gas velocity is much smaller than the speed of light. In Section II.2, we describe the method by Rosdahl and Teyssier (2015), which modifies the M1 closure scheme to follow RT accurately in highly optically thick regions. In Section II.3, we review the RSLA formulation, which relaxes the CFL constraint imposed by the speed of light and allows for larger time steps.
II.1 Basic RHD equations with M1 closure
@@
Here, we introduce the RHD equations. Throughout this paper, we restrict our discussion to the regime where $\beta=|\bm{v}|/c\ll 1$, with $\bm{v}$ denoting the gas velocity and $c$ the speed of light. We take into account both absorption and scattering processes, including their frequency dependence. The formulation of the RHD equations depends on the choice of reference frame, specifically, whether physical quantities are expressed in the lab frame or in the comoving frame. In this study, we adopt the so-called mixed frame formulation, in which opacities and emissivities are evaluated in the comoving frame, while all other gas and radiation quantities are defined in the lab frame (Mihalas and Klein, 1982; Krumholz et al., 2007a). This formulation enables the numerical algorithm to conserve total energy to machine precision during gas–radiation interaction. Moreover, because radiation energy in the lab frame is a conserved quantity, it is obvious how to refine or coarsen the grid in a conservative manner in AMR codes. In this paper, physical quantities measured in the comoving frame are denoted with the subscript 0. We show the Lorentz transformations between the lab frame and comoving frame quantities in Appendix A.
@@
$\kappa_{0E}=\frac{\int_{0}^{\infty}d\nu\hskip 2.84526pt\kappa^{\nu}_{0}E^{\nu}_{\mathrm{rad},0}}{\int_{0}^{\infty}d\nu\hskip 2.84526ptE^{\nu}_{\mathrm{rad},0}},$ (9)
$\chi_{0F}=\frac{\int_{0}^{\infty}d\nu\hskip 2.84526pt(\kappa^{\nu}_{0}+\sigma^{\nu}_{0})F^{\nu}_{0}}{\int_{0}^{\infty}d\nu\hskip 2.84526ptF^{\nu}_{0}},$ (10)
where $\kappa^{\nu}_{0}$ and $\sigma^{\nu}_{0}$ are the frequency-dependent absorption and scattering opacities evaluated in the comoving frame, $B^{\nu}(T)$ is the Planck distribution at temperature $T$, and $E^{\nu}_{\mathrm{rad},0}$ and $F^{\nu}_{0}$ are the comoving frame radiation energy density and flux per unit frequency. For calculating $\kappa_{0E}$ and $\chi_{0F}$ with Equations (9) and (10), we need the spectrum of radiation energy $E^{\nu}_{\mathrm{rad},0}$ and flux $F^{\nu}_{0}$. We describe how to calculate this spectrum with finite frequency bins and low computational cost in Section IV.3.
In the M1 closure method, we determine the radiation pressure tensor $\bm{\mathsf{P}}_{\mathrm{rad}}$ by imposing the M1 closure relation. Specifically, $\bm{\mathsf{P}}_{\mathrm{rad}}$ is expressed in terms of the Eddington tensor $\bm{\mathsf{D}}$ as follows (Levermore, 1984):
@@
With these definitions, the M1 closure reproduces the correct physical behavior of radiation in both the optically thin and optically thick limits. In the optically thin limit where $|\bm{F}|/(cE_{\mathrm{rad}})\simeq 1$ (or $\xi\simeq 1$), the Eddington tensor $\bm{\mathsf{D}}$ approaches $\bm{n}\otimes\bm{n}$, corresponding to a beam-like radiation field. In the optically thick limit where $|F|/(cE_{\mathrm{rad}})\simeq 0$ (or $\xi\simeq 1/3$), it approaches $\bm{\mathsf{I}}/3$, consistent with an isotropic radiation field.
II.2 Streaming/trapped component for an optically thick medium